OpenStudy (anonymous):

How do I solve an exponential equation like this: 37 = 22 + 1(2.2)^3.4x

OpenStudy (anonymous):

take it easy...,

OpenStudy (anonymous):

similar with a = b + c(d)^3 . 4x a - b = c(d)^3 . 4x 4x = (a-b) / c(d)^3 . x = (a-b)/ (4. c(d)^3) i believe you can.., ok

OpenStudy (anonymous):

I don't understand how the third line broke apart the 3.4 in the exponent. I just noticed there's an equation thingy for replies, so I'll use that for the equation just in case. Thank you for replying! \[37 = 22 + 1(2.2)^{3.4x}\]

OpenStudy (anonymous):

I believe it will be algebraically simplified to the following, but I don't know what to do from there: \[15 = 2.2^{3.4x}\]

OpenStudy (anonymous):

yes @BloopityBloop .., maybe using logaritm. \[15 = 2.2^{3.4x}\] \[^{2.2} \log 15 = 3.4x\] so x = i believe you can @BloopityBloop

OpenStudy (anonymous):

\[^{2} \log 8 = 3 \] equal to \[2^{3} = 8\]

OpenStudy (anonymous):

sorry in indonesia.., \[^{a} \log{b} = c \] equal to \[\log_{a}b = c \] how?? can I hep you again?

OpenStudy (anonymous):

yeah, it was the log I'm not too good with. I think I need to take the log of both sides, then simplify with algebra. I got: \[x = \frac{ \ln(15) }{ 3.4 \ln (2.2) }\] Thanks for your help!

OpenStudy (anonymous):

great job @BloopityBloop your the best :)